A Banach space which isn’t metrizable in weak topology.

******************************

Every Hilbert space has this property (cf. [Hal, problem 21]).

Comment. It is probably true that no infinite dimensional Banach space is metrizable in the weak topology. Ref.

[Hal] P.R. Halmos, A Hilbert space problem book, Princeton, Van Nostrand, 1967.